OpenStudy (anonymous):

One of the number of the following is wrong which one is it? 9 10 22 69 281 1405 8436

OpenStudy (zzr0ck3r):

what class is this for?

OpenStudy (anonymous):

its a sat question

OpenStudy (anonymous):

is it 281?

OpenStudy (anonymous):

reason?

OpenStudy (zzr0ck3r):

give the logic @surjithayer

OpenStudy (anonymous):

what about 1405 and 8436?

OpenStudy (anonymous):

9*1+1=10 10*2+2=22 22*3+3=69 69*4+4=28 280*5+5=1405 1405*6+6=8436 an-1*n+n=an

OpenStudy (anonymous):

69*4+4=280

OpenStudy (anonymous):

so the answer is 281 and u rock man